# Information Theory for Complex Systems Scientists: What, Why, & How?

### v3

@@ V.1 Functional Connectivity Networks
The simplest form of network construction is that of a “functional connectivity” (FC) network. In an FC network, the edges between nodes are undirected and describe some symmetric statistical relationship between the two elements of a system. Functional connectivity analysis is generally most well known as a tool in computational neuroscience [154, 164] and an Internet search for “functional connectivity” will return overwhelmingly neuroscience-focused results, however the notion of using time-series statistics to infer a network is much more broadly applicable and has been used when constructing genomic networks [165, 166], financial networks [167, 168], climate networks [169], and developmental biology [76].
@@ V.1.1 Bias in Naive Entropy Estimators
The most common way to estimate entropies and mutual information from discrete data is through a naive estimator, where $P(X=x)$ is a function of how many times $x$ appears in a sample, divided by the total number of samples. Due to finite size effects, however, the naive entropy $\hat{H}(X)$ systematically under-estimates the true entropy, and by extension, over-estimates the true mutual information. Being strictly positive, the naive mutual information estimator is known to be upwardly-biased, reliably over-estimating the true mutual information [84]. This means that, unless large amounts of data are included in the mutual information estimates that define the edge weights in the functional connectivity network, the overall information-density of the system is being over-estimated.
@@ V.1.2 Significance Testing & Thresholding
The other concern when inferring functional connectivity networks is whether an edge is “real”, and deciding on a criteria for whether to include an edge in the network model. The most common way to do this is to compute an analytic or bootstrap a p-value for a null-hypothesis significance test. For Gaussian and discrete variables, it is possible to compute analytic p-values based on Chi-squared distributions [174, 175]. The estimators are infrequently used, although are implemented in the JIDT package [176]. In the absence of an analytic p-value, null hypothesis significance testing can be done by constructing a null distribution (as discussed above and described in detail below). Given some suitably large number of nulls, it is easy to compute an empirical p-value by quantifying the number of null mutual information values greater than or equal to the empirical value.
@@ V.2 Effective Connectivity Networks
The functional connectivity network has no notion of “time” - it describes the instantaneous correlation between elements and would be the same if you were to permute the pair of time series (so long as the joint state was preserved). To account for temporal dynamics requires an alternative approach, one that can infer directed edges. Historically, this was done using the temporal mutual information, quantifying how much knowing the past of $X$ reduces our uncertainty about the future of $Y$. The temporal mutual information has a significant limitation, however: it will conflate true information transfer with active information storage if the signals are autocorrelated. Consider the case of two time series: $X=\sin(t)$ and $Y=\cos(t)$. Since these signals are period and offset by $\pi/2$, knowing the state of $X$ at time $t$ uniquely specifies the state of $Y$ at time $t+\tau$, which is indicated by the non-zero temporal information. However, the same information about $Y_{t+\tau}$ could also be learned by observing $Y$’s own past. There is no additional information disclosed by $X$’s past about $Y$’s future that isn’t disclosed by $Y$’s own past.
@@ V.2.1 Optimizing Parameters
Unlike the functional connectivity, which is parameter-free, the transfer entropy requires deciding on how much of the history from the source and target to consider when computing the conditional mutual information. In Eq. 50, these are the two parameters $k$ (for the source past) and $l$ (for the target past). The question of how best to optimize them is a complex topic, and we will not be able to do it justice here. For readers interested in the gory details, we recommend Bossomaier et al.’s Introduction to Transfer Entropy [84]. Here, we will briefly review the general concerns (over/underestimating the true value), as well as describing a simple, serial conditioning algorithm (based on [106, 187]) and briefly discussing null-models.
@@
Picking the correct embedding then, requires balancing a series of trade-offs. Too little history accounted for will underestimate the true value, while too much history accounted for can compromise it through finite size effects. One possible solution to this problem is a serial conditioning approach (described in detain in [187]). Briefly, the algorithm works by testing whether each additional bin of history produces a statistically significant increase in the transfer entropy when added. For example, if optimizing the target $Y$’s past, one begins with the active information storage with one bin of history: $I(Y_{t-1};Y_{t})$. Then, one bin of history is added in the form of a conditional mutual information: $I(Y_{t-2};Y_{t}|Y_{t-1})$, and this value is significance-tested agianst a null distribution. If the increase in the active information storage is significantly greater than would be expected by chance, then the increase is likely to be real and not simply a finite-size effect. This process can be repeated with increasing values of $k$ (conditioning on all previous $k-1$ values each time) until either subsequent values are no longer significant, or a pre-defined $k_{\max}$ is reached. Once the optimal value $k^{*}$ is selected, the same algorithm can be used to optimize the source dimensionality ($l^{*}$) using the optimal $k^{*}$. It is also possible to use the same logic to infer a non-uniform embedding, given some alternative stopping criterion (such as a pre-selected $k_{\max})$ [189].
@@ V.2.3 Redundancy & Synergy in Effective Connectivity Networks
One possible solution to this is with the global transfer entropy, which conditions the bivariate transfer entropy on the state of every other element in the system. If $X$ and $Y$ are both elements of W:
$TE_{G}(X\rightarrow Y|\textbf{W})=I(X_{-k:t-1};Y_{t}|Y_{-l:t-1},\textbf{W}^{-\{X,Y\}}_{-q:t-1}).$ (89)
Where $\textbf{W}^{-\{X,Y\}}_{-q:t-1}$ indicates the past state of every element of W excluding $X$ and $Y$. The global transfer entropy is conceptually very powerful, as all synergiestic dependencies that inform on the relationship $X\to Y$ will be illuminated, and redundant information will be pruned. This can make it useful for the study of synergies in complex systems (as in [192]). However, for large systems, the amount of data required to estimate the multidimensional probability distribution can be prohibitively large.
To address this, [106] propose a multivariate transfer entropy algorithm that uses robust, multilevel statistical tests and a greedy optimization procedure to identify the minimal set of parent elements $\textbf{W}^{*}$ that optimizes the conditional transfer entropy $TE(X\to Y|\textbf{W}^{*})$. In doing so, the mutlivarite transfer entropy algorithm builds a bivariate network, but where the weights of the edge between source $X$ and target $Y$ is “informed by” the interaction between every other parent of $Y$ as well. Analysis with simulated Gaussian and mean-field field models has found that the multivariate transfer entropy does a much better job at extracting the ground truth generating model of the simulated data then the bivariate transfer entropy [187, 193]. The multivariate transfer entropy network inference algorithm has been applied to inferring interaction networks between cryptocurrencies [91], biological neuronal networks [89, 194], and human electrophysiological data [195].
It is my recommendation that the multivariate transfer entropy analysis should be used whenever plausible (the only existing implementation is provided by the IDTxl package [107]), but it’s limitations should also be made clear. While it is less data-intensive than the global transfer entropy, the optimization of $\textbf{W}^{*}$, like the optimization of the embedding dimensions and lags, requires considerable computational effort in the form of multiple sequences of large null distributions.
@@ V.3 Hypergraphs, Simplicial Complexes, & Other Higher-Order Frameworks
While the multivariate transfer entropy is sensitive to synergies (and prunes redundancies) in a way that the bivariate transfer entropy is not, it could still be argued that networks are simply too restrictive a model for complex systems rich in higher-order redundancies and synergies. For example, consider our toy logical-XOR system (for simplicity, we will exclude the lags from consideration): $Z=X\bigoplus Y$. If we take a conditioning approach, we will find that $I(X;Z|Y)=1$ bit and and $I(Y;Z|X)=1$ bit, which will appear as two, bivariate links in the final network. Unfortunately, looking at the network, it is impossible to differentiate between the case where two links represent a synergistic interaction between two parents and a target or the case where both parents have independent, specific relationships with the target.
@@ VI Integration, Segregation, & Complexity
A core feature of complex systems is that they incorporate elements of “integration” and “segregation” [55]. Integration refers to a dynamic where all of the elements of the system are interacting and affecting each-other, while segregation refers to a dynamic where subsets of elements are involved in their own processes that are not shared with other elements. As an example, consider the brain: it is known that particular brain areas are involved in some processes and not others (functionality is segregated between different regions), however, at the same time, the brain is integrated enough that all the different local processes can be combined into an integrated, unified whole: a unitary organism with an (apparently) single consciousness. It has been hypothesized that this balance of integration and segregation is key for healthy brain function [203, 154]. Similarly, in economics, successful firms maintain a healthy balance of segregation (each branch working on it’s own mission), while all the work is overseen and broadly directed by a centralized executive office. In global politics, the internal dynamics of individual nations are segregated from each-other by national borders, languages, and cultures, while between-nation integration exists in the form of treaties, trade, and historical entanglements.
This mixture of integration and segregation is an inherently multi-scale phenomenon, with different scales often displaying different biases one way or another. Consider a modular network: within each module, there is high integration (which could be assessed using total correlation), but each module may only be sparely connected to other modules, indicating system-wide higher-scale segregation. Depending on how the network is wired together (for instance, having a small-world [204] or scale-free [205] topology), the structure of the overall dynamics being run on the network will change. As mentioned in Sec. I, multi-scale or scale-free structure is one of the hallmarks of complex systems, and is deeply related to the issue of balancing integration and segregation: a system that is fully integrated at the micro-scale cannot support a “higher” scale (a fully connected network has no higher-order communities), nor can a fully segregated system (a set of non-interacting elements barely qualifies as a system). It is a balance of the two that allows non-trivial, multi-scale structure to emerge.
@@ VI.1 TSE-Complexity
$C(\textbf{X})=TC(\textbf{X})-\frac{TC(\textbf{X})}{N}-\mathbb{E}[TC(\textbf{X}_{-i})]$ (92)
@@ VI.2 O-Information & S-Information
The O-information (which was originally named the “enigmatic information” by James and Crutchfield [214]) has the useful property that if X is dominated by synergistic interactions, then $\Omega({\textbf{X}})<0$ bit and conversely, if X is dominated by redundant interactions, $\Omega({\textbf{X}})>0$ bit. It is 0 bit if X is completely disintegrated (i.e. all $X_{i}$ are independent), and it is only sensitive to beyond-pairwise dependencies (i.e. the bivariate mutual information doesn’t contribute towards $+\Omega$ or $-\Omega$). A negative value of $\Omega(\textbf{X})$ only indicates that X is dominated by synergistic dependencies - redundant dependencies can also exist, which makes it more of a heuristic than the complete decompositions provided by the PID and PED. The O-information has generated considerable interest in diverse fields, including neuroscience [215, 211], psychology [marinazzo_information-theoretic_2022], and music theory [216]. Unlike the PID, which grows hyperexponentially with the number of elements in the system, the O-information scales much more gracefully and can be computed for several hundred interacting elements (assuming sufficient data). The measure has also been extended in a number of directions, including a modified transfer entropy [217], and a time-lagged, frequency-specific version [218].
@@
$\Omega(\textbf{X})=(2-N)TC(\textbf{X})+\sum_{i=1}^{N}TC(\textbf{X}_{-i})$ (95)
This can help provide some intuition into what synergy means. The left-side term $(2-N)TC(\textbf{X})$ is the total integration of the “whole”, duplicated $N-2$ times (and made negative). The right-hand term “adds back” the total correlation associated with the $N$ possible ways that single elements can be removed. If $\sum_{i=1}^{N}TC(\textbf{X}_{-i}>(2-N)TC(\textbf{X})$, then there must be correlations in the $\textbf{X}_{-i}$ that are redundant, and so get double counted. Conversely, if $(2-N)TC(\textbf{X})$ is greater, then there is information in the structure of the “whole” that can’t be found in the parts when all added together.
The O-information can be analyzed locally, in the same way that the bivariate mutual information can be [216]. For a particular configuration x for X:
@@
Where $tc$ and $dtc$ are the local total correlation and dual total correlation (which can be computed from local entropies). Generally, the sign of $\omega$ is interpreted in the same way as the in the expected case: a negative sign indicates greater synergy while a positive sign indicates greater redundancy. The local total correlation and local dual total correlation can both be negative, however (unlike their strictly non-negative expected forms). This can create a situation: suppose that $tc(\textbf{x})=-3.0$ bit and $dtc(\textbf{x})=-8.0$ bit. In this case, $\omega(\textbf{x})=+5.0$ bit and we must ask: if $dtc$ is more misinformative than the $tc$, is that the same thing as the overall system being “redundancy dominated”? The exploration of the local O-information at this point is very limited and so there remain fundamental questions waiting to be answered.
@@ VI.3 Whole-Minus-Sum Complexity & Integrated Information
$\Phi^{R}(\textbf{X})=\Phi^{WMS}(\textbf{X})-I_{\partial\cap}(\{1\}\{2\}\to\{1\}\{2\})$ (102)
@@
Despite these limitations however, we are of the opinion that “weak integrated information theory” holds tremendous promise that has only begun to be realized. It provides an appealing link between the worlds of information theory and computation and dynamical systems approaches on the other. It also is built on strong axiomatic foundations (Shannon information theory, and extensions), which make measures interpretable in ways that other complexity measures may not be. And finally, being purely based on information theory, it is extremely general and can be applied to a variety of domains.

### v4

@@ V.1 Functional Connectivity Networks
The simplest form of network construction is that of a “functional connectivity” (FC) network. In an FC network, the edges between nodes are undirected and describe some symmetric statistical relationship between the two elements of a system. Functional connectivity analysis is generally most well known as a tool in computational neuroscience [156, 166] and an Internet search for “functional connectivity” will return overwhelmingly neuroscience-focused results, however the notion of using time-series statistics to infer a network is much more broadly applicable and has been used when constructing genomic networks [167, 168], financial networks [169, 170], climate networks [171], and in developmental biology [172].
@@ V.1.1 Bias in Naive Entropy Estimators
The most common way to estimate entropies and mutual information from discrete data is through a naive estimator, where $P(X=x)$ is a function of how many times $x$ appears in a sample, divided by the total number of samples. Due to finite size effects, however, the naive entropy $\hat{H}(X)$ systematically under-estimates the true entropy, and by extension, over-estimates the true mutual information. Being strictly positive, the naive mutual information estimator is known to be upwardly-biased, reliably over-estimating the true mutual information [83]. This means that, unless large amounts of data are included in the estimates that define the edge weights, the overall information-density of the system is being over-estimated.
@@ V.1.2 Significance Testing & Thresholding
The other concern when inferring functional connectivity networks is whether an edge is “real”, and deciding on a criteria for whether to include an edge in the network model. The most common way to do this is to compute an analytic or bootstraped p-value for a null-hypothesis significance test. For Gaussian and discrete variables, it is possible to compute analytic p-values based on Chi-squared distributions [177, 178]. The estimators are infrequently used, although are implemented in the JIDT package [179]. In the absence of an analytic p-value, null hypothesis significance testing can be done by constructing a null distribution (as discussed above and described in detail below). Given some suitably large number of nulls, it is easy to compute an empirical p-value by quantifying the number of null mutual information values greater than or equal to the empirical value.
@@ V.2 Effective Connectivity Networks
The functional connectivity network has no notion of “time” - it describes the instantaneous correlation between elements and would be the same if one were to permute frames (preserving the joint state). To account for temporal dynamics requires an alternative approach, one that can infer directed edges. Historically, this was done using the temporal mutual information, quantifying how much knowing the past of $X$ reduces our uncertainty about the future of $Y$. The temporal mutual information has a significant limitation, however: it will conflate true information transfer with active information storage if the signals are autocorrelated. Consider the case of two time series: $X=\sin(t)$ and $Y=\cos(t)$. Since these signals are periodic and offset by $\pi/2$, knowing the state of $X$ at time $t$ uniquely specifies the state of $Y$ at time $t+\tau$, which is indicated by the non-zero temporal information. However, the same information about $Y_{t+\tau}$ could also be learned by observing $Y$’s own past. There is no additional information disclosed by $X$’s past about $Y$’s future that isn’t disclosed by $Y$’s own past.
@@ V.2.1 Optimizing Parameters
Unlike the functional connectivity, which is parameter-free, the transfer entropy requires deciding on how much of the history from the source and target to consider when computing the conditional mutual information. In Eq. 50, these are the two parameters $k$ (for the source past) and $l$ (for the target past). The question of how best to optimize them is a complex topic, and I will not be able to do it justice here. For readers interested in the gory details, I recommend Bossomaier et al.’s Introduction to Transfer Entropy [83]. Here, I will briefly review the general concerns (over/underestimating the true value), as well as describing a simple, serial conditioning algorithm (based on [106, 190]) and briefly discussing null-models (again).
@@
Picking the correct embedding then, requires balancing a series of trade-offs. Too little history accounted for will underestimate the true value, while too much history accounted for can compromise it through finite size effects. One possible solution to this problem is a serial conditioning approach (described in detail by [190]). Briefly, the algorithm works by testing whether each additional bin of history produces a statistically significant increase in the transfer entropy when added. For example, if optimizing the target $Y$’s past, one begins with the active information storage with one bin of history: $I(Y_{t-1};Y_{t})$. Then, one bin of history is added in the form of a conditional mutual information: $I(Y_{t-2};Y_{t}|Y_{t-1})$, and this value is significance-tested against a null distribution. If the increase in the active information storage is significantly greater than would be expected by chance, then the increase is likely to be real and not simply a finite-size effect. This process can be repeated with increasing values of $k$ (conditioning on all previous $k-1$ values each time) until either subsequent values are no longer significant, or a pre-defined $k_{\max}$ is reached. Once the optimal value $k^{*}$ is selected, the same algorithm can be used to optimize the source dimensionality ($l^{*}$) using the optimal $k^{*}$. It is also possible to use the same logic to infer a non-uniform embedding, given some alternative stopping criterion (such as a pre-selected $k_{\max})$ [192].
@@ V.2.3 Redundancy & Synergy in Effective Connectivity Networks
One possible solution to this is with the global transfer entropy, which conditions the bivariate transfer entropy on the state of every other element in the system. If $X$ and $Y$ are both elements of V:
$TE_{G}(X\rightarrow Y|\textbf{V})=I(X_{-k:t-1};Y_{t}|Y_{-l:t-1},\textbf{V}^{-\{X,Y\}}_{-q:t-1}).$ (89)
Where $\textbf{V}^{-\{X,Y\}}_{-q:t-1}$ indicates the past state of every element of W excluding $X$ and $Y$. The global transfer entropy is conceptually very powerful, as all synergiestic dependencies that inform on the relationship $X\to Y$ will be illuminated, and redundant information will be pruned. This can make it useful for the study of synergies in complex systems (as in [195]). However, for large systems, the amount of data required to estimate the multidimensional probability distribution can be prohibitively large.
To address this, [106] propose a multivariate transfer entropy algorithm that uses robust, multilevel statistical tests and a greedy optimization procedure to identify the minimal set of parent elements $\textbf{V}^{*}$ that optimizes the conditional transfer entropy $TE(X\to Y|\textbf{V}^{*})$. In doing so, the mutlivarite transfer entropy algorithm builds a bivariate network, but where the weights of the edge between source $X$ and target $Y$ is “informed by” the interaction between every other parent of $Y$ as well. Analysis with simulated Gaussian and mean-field field models has found that the multivariate transfer entropy does a much better job at extracting the ground truth generating model of the simulated data then the bivariate transfer entropy [190, 196]. The multivariate transfer entropy network inference algorithm has been applied to inferring interaction networks between cryptocurrencies [91], biological neuronal networks [89, 197], and human electrophysiological data [198].
It is my recommendation that the multivariate transfer entropy analysis should be used whenever plausible (the only existing implementation is provided by the IDTxl package [107]), but its limitations should also be made clear. While it is less data-intensive than the global transfer entropy, the optimization of $\textbf{V}^{*}$, like the optimization of the embedding dimensions and lags, requires considerable computational effort in the form of multiple sequences of large null distributions. This can produce extremely long runtimes as multiple optimizations are run in sequence.
@@ V.3 Hypergraphs, Simplicial Complexes, & Other Higher-Order Frameworks
While the multivariate transfer entropy is sensitive to synergies (and prunes redundancies) in a way that the bivariate transfer entropy is not, it could still be argued that networks are simply too restrictive a model for complex systems rich in higher-order redundancies and synergies. For example, consider our toy logical-XOR system (for simplicity, we will exclude the lags from consideration): $Z=X\bigoplus Y$. If we take a conditioning approach, we will find that $I(X;Z|Y)=1$ bit and and $I(Y;Z|X)=1$ bit, which will appear as two, bivariate links in the final network. Unfortunately, looking at the network, it is impossible to differentiate between the case where two links represent a synergistic interaction between two parents and a target versus the case where both parents have independent, specific relationships with the target.
@@ VI Integration, Segregation, & Complexity
A core feature of complex systems is that they incorporate elements of “integration” and “segregation” [56]. Integration refers to a dynamic where all of the elements of the system are interacting and affecting each-other, while segregation refers to a dynamic where subsets of elements are involved in their own processes that are not shared with other elements. As an example, consider the brain: it is known that particular brain areas are involved in some processes and not others (functionality is segregated between different regions), however, at the same time, the brain is integrated enough that all the different local processes can be combined into an integrated, unified whole: a unitary organism with an (apparently) single consciousness. It has been hypothesized that this balance of integration and segregation is key for healthy brain function [206, 156]. Similarly, in economics, successful firms maintain a healthy balance of segregation (each branch working on its own mission), while all the work is overseen and broadly directed by a centralized executive office. In global politics, the internal dynamics of individual nations are segregated from each-other by national borders, languages, and cultures, while between-nation integration exists in the form of treaties, trade, and historical entanglements.
This mixture of integration and segregation is an inherently multi-scale phenomenon, with different scales often displaying different biases one way or another. Consider a modular network: within each module, there is high integration (which could be assessed using total correlation [207]), but each module may only be sparely connected to other modules, indicating system-wide higher-scale segregation. Depending on how the network is wired together (for instance, having a small-world [208] or scale-free [209] topology), the structure of the overall dynamics being run on the network will change. As mentioned in Sec. I, multi-scale or scale-free structure is one of the hallmarks of complex systems, and is deeply related to the issue of balancing integration and segregation: a system that is fully integrated at the micro-scale cannot support a “higher” scale (a fully connected network has no higher-order communities), nor can a fully segregated system (a set of non-interacting elements barely qualifies as a system). It is a balance of the two that allows non-trivial, multi-scale structure to emerge.
@@ VI.1 TSE-Complexity
$C(\textbf{X})=TC(\textbf{X})-\frac{TC(\textbf{X})}{N}-\mathbb{E}[TC(\textbf{X}^{-i})]$ (92)
@@ VI.2 O-Information & S-Information
The O-information (which was originally named the “enigmatic information” by James and Crutchfield [218]) has the useful property that if X is dominated by synergistic interactions, then $\Omega({\textbf{X}})<0$ bit and conversely, if X is dominated by redundant interactions, $\Omega({\textbf{X}})>0$ bit. It is 0 bit if X is completely disintegrated (i.e. all $X_{i}$ are independent), and it is only sensitive to beyond-pairwise dependencies (i.e. the bivariate mutual information doesn’t contribute towards $+\Omega$ or $-\Omega$). A negative value of $\Omega(\textbf{X})$ only indicates that X is dominated by synergistic dependencies - redundant dependencies can also exist, which makes it more of a heuristic than the complete decompositions provided by the PID and PED. The O-information has generated considerable interest in diverse fields, including neuroscience [219, 215], psychology [203], and music theory [220]. Unlike the PID, which grows hyperexponentially with the number of elements in the system, the O-information scales much more gracefully and can be computed for several hundred interacting elements (assuming sufficient data). The measure has also been extended in a number of directions, including a modified transfer entropy [221], a local perspective to capture time-resolved redundancy/synergy dynamics [220, 48], and a time-lagged, frequency-specific version [222].
@@
$\Omega(\textbf{X})=(2-N)TC(\textbf{X})+\sum_{i=1}^{N}TC(\textbf{X}^{-i})$ (95)
This can help provide some intuition into what synergy means. The left-side term $(2-N)TC(\textbf{X})$ is the total integration of the “whole”, duplicated $N-2$ times (and made negative). The right-hand term “adds back” the total correlation associated with the $N$ possible ways that single elements can be removed. If $\sum_{i=1}^{N}TC(\textbf{X}^{-i})>(2-N)TC(\textbf{X})$, then there must be correlations in the $\textbf{X}^{-i}$ that are redundant, and so get double counted. Conversely, if $(2-N)TC(\textbf{X})$ is greater, then there is information in the structure of the “whole” that can’t be found in the parts when all added together.
The O-information can be analyzed locally, in the same way that the bivariate mutual information can be [220, 48]. For a particular configuration x of X:
@@
Where $tc$ and $dtc$ are the local total correlation and dual total correlation (which can be computed from local entropies). Generally, the sign of $\omega$ is interpreted in the same way as the in the expected case: a negative sign indicates greater synergy while a positive sign indicates greater redundancy. The local total correlation and local dual total correlation can both be negative, however (unlike their strictly non-negative expected forms). This can create more complex situations: suppose that $tc(\textbf{x})=-3.0$ bit and $dtc(\textbf{x})=-8.0$ bit. In this case, $\omega(\textbf{x})=+5.0$ bit and we must ask: if $dtc$ is more misinformative than the $tc$, is that the same thing as the overall system being “redundancy dominated”? The exploration of the local O-information at this point is very limited (although see [48] for recent work) and so there remain fundamental questions waiting to be answered.
@@ VI.3 Whole-Minus-Sum Complexity & Integrated Information
$\Phi^{R}(\textbf{X})=\Phi^{WMS}(\textbf{X})-I_{\partial}(\{1\}\{2\}\to\{1\}\{2\})$ (102)
@@
Despite these limitations however, I am of the opinion that “weak integrated information theory” holds tremendous promise that has only begun to be realized. It provides an appealing link between the worlds of information theory and computation and dynamical systems approaches on the other. It also is built on strong axiomatic foundations (Shannon information theory, and extensions), which make measures interpretable in ways that other complexity measures may not be. And finally, being purely based on information theory, it is extremely general and can be applied to a variety of domains.